The complex that is formed between x and y has an absorption maximum of 361nm. A 0.0075M solution of the complex has a transmittance of 29.8% and measured in a 1.00 cm cell at this wavelength. Calculate the molar absorbency of the complex
Answer #1

wavelength =361nm

length of the cuvette l=1 cm

concentration of solution c= 0.0075 M

let the molar absorptivity of the complex be

from lambert Beer law

absorbance A =cl equation 1

if I0 is the intensity of light and I is the intensity leaving the sample

A=log(I0/I)

and transmittance is defined as

T= log(I/I0)

T%=log(I/I0)*100

there we see

A=log(1/T)

A=log(100/%T)

A=2-log(%T)

given T%=29.8 %

just plug in the values to get value of Absorbance

A=2-log(%T)

A=2-log(29.8)

A= 2-1.474

A=0.526

using equation 1

A =cl

0.526= *1cm*  0.0075M

0.526/(1cm*  0.0075M)=

70.133 cm-1M-1= (Answer)